# Decoupled Data Consistency with Diffusion Purification for Image Restoration

### v5

@@ A.1 Discussion on Hyperparameters
A.2 Choice of Hyperparameters
In this section, we specify the hyperparameters used for generating the results in Section 5. For the baselines, we either followed the instructions given in each respective work or used the parameters provided in their code.
A.2.1 Experiments with Pixel Space Diffusion Models
For the hyperparameters used for the pixel space experiments, we tabulate all of our values for each dataset in Table 6. Recall that for $T_{k}$, we choose a linearly decaying time schedule. We start at a value $T_{0}$, which is specified in the table. For box inpainting, we used a larger value of $T_{0}$ to compensate for the large artifacts that exist within the box region. The diffusion purification is performed with 20 deterministic DDIM steps.
We optimize the data fidelity objective with momentum-based gradient descent with the momentum hyperparameter set to $0.9$ for both FFHQ and LSUN-Bedrooms datasets. To achieve the best performance, we use Pixel Space Optimization for Gaussian deblurring and motion deblurring while using Latent Space Optimization for box inpainting and super-resolution. Again, we fix the total number of gradient steps as 1000 throughout the $K$ iterations, which means we take $\frac{1000}{K}$ gradient steps in each round of data fidelity optimization.
@@ A.2.1 Experiments with Pixel Space Diffusion Models
$\alpha$ $K$ $T_{0}$ $\alpha$ $K$ $T_{0}$ $\alpha$ $K$ $T_{0}$ $\alpha$ $K$ $T_{0}$
LSUN-Bedrooms $10^{3}$ $10$ $400$ $10^{5}$ $10$ $400$ $10^{5}$ $20$ $400$ $10^{3}$ $20$ $700$
FFHQ $10^{3}$ 10 $400$ $10^{5}$ $10$ $400$ $10^{5}$ $20$ $400$ $10^{3}$ $20$ $700$
Table 6: Table of hyperparameters used for experiments in the pixel space. $\alpha$ refers to the learning rate used in the data consistency optimization, $K$ refers to the total number of iterations, and $T_{0}$ refers to the starting point for the linearly decaying time schedule.
@@
$\alpha$ $K$ $T_{0}$ $\alpha$ $K$ $T_{0}$ $\alpha$ $K$ $T_{0}$ $\alpha$ $K$ $T_{0}$
LSUN-Bedrooms $10^{3}$ $10$ $400$ $10^{5}$ $10$ $400$ $10^{5}$ $10$ $400$ $10^{3}$ $20$ $500$
FFHQ $10^{3}$ $10$ $400$ $10^{5}$ $10$ $400$ $10^{5}$ $4$ $400$ $10^{3}$ $20$ $500$
Table 7: Table of hyperparameters used for experiments in the latent space. $\alpha$ refers to the learning rate used in the data consistency optimization, $K$ refers to the total number of iterations, and $T_{0}$ refers to the starting point for the linearly decaying time schedule.
@@
$\alpha$ $K$ $T_{0}$ $\alpha$ $K$ $T_{0}$ $\alpha$ $K$ $T_{0}$ $\alpha$ $K$ $T_{0}$
LSUN-Bedrooms $10^{3}$ $20$ $1$ $10^{5}$ $20$ $1$ $10^{5}$ $20$ $1$ $10^{3}$ $20$ $5$
Table 8: Table of hyperparameters used for experiments with CMs. $\alpha$ refers to the learning rate used in the data consistency optimization, $K$ refers to the total number of iterations, and $T_{0}$ refers to the starting point for the linearly decaying time schedule. Different from the typical diffusion models used in previous experiments which use a noise scale in the range of [0,1000], CMs are trained in the range of [0.002,80].
@@
Ours Pixel (Version 1) $\bm{0.154}\pm 0.03$ $\underline{29.43}\pm 2.01$ $\underline{0.844}\pm 0.03$ $\underline{0.052}\pm 0.01$ $\underline{25.54}\pm 3.01$ $\underline{0.917}\pm 0.01$
Ours Pixel (Version 2) $0.185\pm 0.03$ $\bm{30.86}\pm 2.11$ $\bm{0.878}\pm 0.03$ $\bm{0.051}\pm 0.01$ $\bm{28.02}\pm 2.74$ $\bm{0.935}\pm 0.01$
@@
Ours LDMs (Version 1) $\bm{0.183}\pm 0.04$ $\underline{29.16}\pm 1.65$ $\underline{0.826}\pm 0.03$ $\bm{0.166}\pm 0.03$ $\underline{26.50}\pm 2.20$ $\underline{0.860}\pm 0.02$
Ours LDMs (Version 2) $\underline{0.215}\pm 0.04$ $\bm{30.61}\pm 1.97$ $\bm{0.866}\pm 0.03$ $\underline{0.204}\pm 0.04$ $\bm{27.47}\pm 2.05$ $\bm{0.868}\pm 0.03$
@@
Ours Pixel (Version 1) $\bm{0.140}\pm 0.02$ $\underline{30.41}\pm 1.93$ $\underline{0.864}\pm 0.03$ $\bm{0.051}\pm 0.01$ $\underline{37.56}\pm 1.77$ $\underline{0.963}\pm 0.01$
Ours Pixel (Version 2) $0.173\pm 0.03$ $\bm{31.67}\pm 2.12$ $\bm{0.891}\pm 0.03$ $\underline{0.060}\pm 0.01$ $\bm{38.37}\pm 1.88$ $\bm{0.970}\pm 0.01$
@@
Ours LDMs (Version 1) $\underline{0.198}\pm 0.03$ $\underline{28.91}\pm 1.98$ $\underline{0.825}\pm 0.04$ $\underline{0.152}\pm 0.03$ $\underline{30.11}\pm 1.93$ $\underline{0.851}\pm 0.03$
Ours LDMs (Version 2) $\bm{0.171}\pm 0.02$ $\bm{30.45}\pm 1.86$ $\bm{0.877}\pm 0.03$ $\bm{0.108}\pm 0.02$ $\bm{31.48}\pm 1.94$ $\bm{0.894}\pm 0.02$
@@
A.2.2 Experiments with LDMs
Similarly, we specify the hyperparameters used for experiments using LDMs in Table 7. For box inpainting, similar to the pixel space experiments, we used a larger value of $T_{0}$ to compensate for the artifacts seen in the box region. The diffusion purification is performed with 20 deterministic DDIM steps. For the data fidelity objectives, the same procedure done in the pixel space was applied for the LDM case.
A.2.3 Experiments with CMs
Lastly, we specify the hyperparameters used for experiments using CMs in Table 8. Notice that different from the typical diffusion model used previously which use a noise scale in the range of [0,1000], CMs are traind in the range of [0.002,80]. Again, the diffusion purification is performed with 20 deterministic DDIM steps.
A.3 Dealing with Measurement Noise
When the measurements are noiseless, a linear decaying schedule starts from $T_{0}$ and ends at $T_{K}=0$ is capable of resulting in high quality reconstructions and we can use a relatively high learning rate for the data fidelity optimization. On the contrary, in the presence of measurement noise, the data fidelity optimization in (10) will inevitably introduce additional artifacts even if $\bm{v}_{k-1}$ is the ground truth image. To achieve robust recovery, one may instead: (i) use a relatively small learning rate for data fidelity optimization to prevent overfitting to the noise, and (ii) end the time schedule at $t_{K}=T$, where $T>0$, to compensate for the artifacts resulting from the measurement noise. As shown in  Figure 16, compared to the clean measurements case, due to the additive noise, the data fidelity optimization always introduce artifacts in $\bm{x}_{k}$. This issue can be alleviated by utilizing a non-zero diffusion purification strength at the final iteration. For the experiments in Section 5.3, all the hyperparameters are kept the same as described in Section A.2.1, but with the following changes:
@@ A.3 Dealing with Measurement Noise
  • The time schedules for super-resolution and motion deblurring start from $t_{0}=400$ and decay linearly to $250$ and $300$ respectively.
@@
Appendix B Additional Quantitative Results
In this section, we provide additional quantitative results for solving inverse problems on the FFHQ dataset. The results are presented in Tables 9 and 10. Throughout all of our experiments, we observe that our proposed algorithms outperform the baselines for various image restoration problems.
(a) Reference and Measurements for Super-Resolution (4x). The noisy measurements are corrupted by random Gaussian noise with $\sigma=0.1.$
(b) Reconstruction Process with Clean Measurements. The reconstruction process takes $K=10$ iterations in total with a time schedule starting at $T_{1}=400$ and linearly decaying to $T_{10}=0$. The learning rate for data fidelity optimization is $lr=1000$. Due to the space limit, only reconstructions at timesteps 1, 3, 5, 7 and 10 are shown.
(c) Reconstruction Process with Noisy Measurements. To compensate the artifacts introduced by the measurement noise, the reconstruction process takes $K=10$ iterations in total with a time schedule starting at $T_{1}=400$ and linearly decaying to $T_{10}=200$. The learning rate for data fidelity optimization is $lr=100$.
Figure 16: Reconstruction Process for Super-Resolution (4x) with Pixel Space Diffusion Models.

### v6

@@ Appendix A Discussion on Hyperparameters
Appendix B Choice of Hyperparameters
In this section, we specify the hyperparameters of DCDP for generating the results in Section 4. For the baselines, we tune the hyperparameters via grid search to obtain the best possible performance.
Experiments with Pixel Space Diffusion Models
For the pixel space experiments, we tabulate the hyperparameters for each dataset in Table 6 (linear inverse problems) and in Table 8 (nonlinear deblurring). Recall that for $t_{k}$, we use a linearly decaying time schedule, which starts at a initial value $t_{1}$ and ends at $t_{K}$. For linear inverse problems, the DCDP-DDIM and DCDP-Tweedie share the same set of hyperparameters. The diffusion purification in DCDP-DDIM is performed with 20 deterministic DDIM steps.
We optimize the data fidelity objective with momentum-based gradient descent with the momentum hyperparameter set to $0.9$ for all tasks. For linear inverse problems, we fix the total number of gradient steps as 1000 throughout the $K$ iterations, which means $\tau=\frac{1000}{K}$ in each round of data fidelity optimization. For nonlinear deblurring, we instead use a total number of 2000 gradients steps, i.e., $\tau=\frac{2000}{K}$ in each round of data fidelity optimization.
@@ Appendix B Choice of Hyperparameters
$lr$ $K$ $\tau$ $t_{1}$ $t_{K}$ $lr$ $K$ $\tau$ $t_{1}$ $t_{K}$ $lr$ $K$ $\tau$ $t_{1}$ $t_{K}$ $lr$ $K$ $\tau$ $t_{1}$ $t_{K}$
LSUN-Bedrooms $10^{3}$ $10$ $100$ $400$ 00 $10^{5}$ $10$ $100$ $400$ 00 $10^{5}$ $20$ $50$ $400$ 00 $10^{3}$ $20$ $50$ $700$ 00
FFHQ $10^{3}$ $10$ $100$ $400$ 00 $10^{5}$ $10$ $100$ $400$ 00 $10^{5}$ $20$ $50$ $400$ 00 $10^{3}$ $20$ $50$ $700$ 00
Table 6: Hyperparameters of DCDP with pixel space diffusion models for solving linear inverse problems. For linear inverse problems, DCDP-Tweedie and DCDP-DDIM share the same set of hyperparameters.
@@
$lr$ $K$ $\tau$ $t_{1}$ $t_{K}$ $lr$ $K$ $\tau$ $t_{1}$ $t_{K}$ $lr$ $K$ $\tau$ $t_{1}$ $t_{K}$ $lr$ $K$ $\tau$ $t_{1}$ $t_{K}$
LSUN-Bedrooms $10^{3}$ $10$ $100$ $400$ 00 $10^{5}$ $10$ $100$ $400$ 00 $10^{5}$ $10$ $100$ $400$ 00 $10^{3}$ $20$ $50$ $500$ 00
FFHQ $10^{3}$ $10$ $100$ $400$ 00 $10^{5}$ $4$ $250$ $400$ 00 $10^{5}$ $10$ $100$ $400$ 00 $10^{3}$ $20$ $50$ $500$ 00
Table 7: Hyperparameters of DCDP with latent space diffusion models for solving linear inverse problems. For linear inverse problems, DCDP-Tweedie and DCDP-DDIM share the same set of hyperparameters.
Dataset DCDP-DDIM DCDP-Tweedie DCDP-LDM-DDIM DCDP-LDM-Tweedie
$lr$ $K$ $\tau$ $t_{1}$ $t_{K}$ $lr$ $K$ $\tau$ $t_{1}$ $t_{K}$ $lr$ $K$ $\tau$ $t_{1}$ $t_{K}$ $lr$ $K$ $\tau$ $t_{1}$ $t_{K}$
ImageNet $10^{3}$ $20$ $100$ $400$ $100$ $10^{3}$ $20$ $100$ $400$ $50$ $5\times 10^{3}$ $20$ $100$ $300$ $50$ $5\times 10^{3}$ $20$ $100$ $300$ $50$
Table 8: Hyperparameters of DCDP for solving nonlinear deblurring.
@@
$lr$ $K$ $\tau$ $t_{1}$ $t_{K}$ $lr$ $K$ $\tau$ $t_{1}$ $t_{K}$ $lr$ $K$ $\tau$ $t_{1}$ $t_{K}$ $lr$ $K$ $\tau$ $t_{1}$ $t_{K}$
LSUN-Bedrooms $10^{3}$ $20$ $50$ $1$ 00 $10^{5}$ $20$ $50$ $1$ 00 $10^{5}$ $20$ $50$ $1$ 00 $10^{3}$ $20$ $50$ $5$ 00
Table 9: Hyperparameters of DCDP with CMs. Different from the diffusion models in previous experiments which specify the noise level via discrete forward steps (ranging from 0 to 1,000), CMs are trained on a continous range of Gaussian noise with standard deviation spanning [0.002,80]. Therefore, $t_{1}$ and $t_{K}$ in this table stands for the standard deviation of the noise rather than the timestep.
@@
DDNM [13] $0.197\pm 0.04$ $28.03\pm 1.46$ $0.795\pm 0.08$ $0.235\pm 0.06$ $24.47\pm 2.38$ $0.837\pm 0.08$
RED-Diff [45] $0.212\pm 0.06$ $27.18\pm 1.23$ $0.781\pm 0.07$ $0.165\pm 0.08$ $24.22\pm 3.16$ $0.824\pm 0.05$
DiffPIR [50] $0.260\pm 0.06$ $26.64\pm 1.78$ $0.767\pm 0.06$ $0.108\pm 0.07$ $23.89\pm 1.89$ $0.814\pm 0.06$
DCDP-DDIM $\bm{0.154}\pm 0.03$ $\underline{29.43}\pm 2.01$ $\underline{0.844}\pm 0.03$ $\underline{0.052}\pm 0.01$ $\underline{25.54}\pm 3.01$ $\underline{0.917}\pm 0.01$
DCDP-Tweedie $0.185\pm 0.03$ $\bm{30.86}\pm 2.11$ $\bm{0.878}\pm 0.03$ $\bm{0.051}\pm 0.01$ $\bm{28.02}\pm 2.74$ $\bm{0.935}\pm 0.01$
@@
DCDP-LDM-DDIM $\bm{0.183}\pm 0.04$ $\underline{29.16}\pm 1.65$ $\underline{0.826}\pm 0.03$ $\bm{0.166}\pm 0.03$ $\underline{26.50}\pm 2.20$ $\underline{0.860}\pm 0.02$
DCDP-LDM-Tweedie $\underline{0.215}\pm 0.04$ $\bm{30.61}\pm 1.97$ $\bm{0.866}\pm 0.03$ $\underline{0.204}\pm 0.04$ $\bm{27.47}\pm 2.05$ $\bm{0.868}\pm 0.03$
@@
DCDP-DDIM $\bm{0.140}\pm 0.02$ $\underline{30.41}\pm 1.93$ $\underline{0.864}\pm 0.03$ $\bm{0.051}\pm 0.01$ $\underline{37.56}\pm 1.77$ $\underline{0.963}\pm 0.01$
DCDP-Tweedie $0.173\pm 0.03$ $\bm{31.67}\pm 2.12$ $\bm{0.891}\pm 0.03$ $\underline{0.060}\pm 0.01$ $\bm{38.37}\pm 1.88$ $\bm{0.970}\pm 0.01$
@@
DCDP-LDM-DDIM $\underline{0.198}\pm 0.03$ $\underline{28.91}\pm 1.98$ $\underline{0.825}\pm 0.04$ $\underline{0.152}\pm 0.03$ $\underline{30.11}\pm 1.93$ $\underline{0.851}\pm 0.03$
DCDP-LDM-Tweedie $\bm{0.171}\pm 0.02$ $\bm{30.45}\pm 1.86$ $\bm{0.877}\pm 0.03$ $\bm{0.108}\pm 0.02$ $\bm{31.48}\pm 1.94$ $\bm{0.894}\pm 0.02$
@@
Experiments with LDMs
We specify the hyperparameters used for experiments using LDMs in Table 7 (linear inverse problems) and in Table 8 (nonlinear deblurring). Again, DCDP-LDM-DDIM and DCDP-LDM-Tweedie share the same set of hyperparameters for linear inverse problems. The diffusion purification in DCDP-LDM-DDIM is performed with 20 deterministic DDIM steps and for the data fidelity objectives, the same procedure done in the pixel space is applied on the LDMs. To achieve the best performance, we use Pixel Space Optimization for Gaussian deblurring and motion deblurring while using Latent Space Optimization for box inpainting, super-resolution and nonlinear deblurring. Again, we fix the total number of gradient steps for linear inverse problems as 1000 while using 2000 steps for nonlinear deblurring.
Experiments with CMs
Lastly, we specify the hyperparameters used for experiments using CMs in Table 9. Note that different from the typical diffusion models which specify the noise level via discrete forward steps (commonly ranging from 0 to 1,000), CMs are trained on a continuous range of Gaussian noise with standard deviation spanning [0.002,80].
Appendix C Dealing with Measurement Noise
When the measurements are noiseless, a linear decaying schedule starting from $t_{1}$ and ending at $t_{K}=0$ is capable of resulting in high-quality reconstructions, and we can use a relatively high learning rate for the data fidelity optimization. On the contrary, in the presence of measurement noise, the data fidelity optimization in (18) will inevitably introduce additional artifacts even if $\bm{v}_{k-1}$ is the ground truth image. To achieve robust recovery, one may instead: (i) use a relatively small learning rate for data fidelity optimization to prevent overfitting to the noise, and (ii) end the time schedule at $t_{K}=\tilde{t}$, where $\tilde{t}>0$, to compensate for the artifacts resulting from the measurement noise. As shown in Figure 19, compared to the clean measurements case, due to the additive noise, the data fidelity optimization always introduces artifacts in $\bm{x}_{k}$. This issue can be alleviated by utilizing a non-zero diffusion purification strength at the final iteration. For the experiments in Section 4.3, all the hyperparameters are kept the same as described in Appendix B, but with the following changes:
@@ Appendix C Dealing with Measurement Noise
  • The time schedules for super-resolution and motion deblurring start from $t_{1}=400$ and decay linearly to $250$ and $300$ respectively.
@@
Figure 17: Reference and Measurements for Super-Resolution (4x). The noisy measurements are corrupted by random Gaussian noise with $\sigma=0.1$.
Figure 18: Reconstruction Process with Clean Measurements. The reconstruction process takes $K=10$ iterations in total with a time schedule starting at $t_{1}=400$ and linearly decaying to $t_{10}=0$. The learning rate for data fidelity optimization is $lr=1000$. Due to the space limit, only reconstructions at timesteps 1, 3, 5, 7 and 10 are shown.
Figure 19: Reconstruction Process with Noisy Measurements. To compensate the artifacts introduced by the measurement noise, the reconstruction process takes $K=10$ iterations in total with a time schedule starting at $t_{1}=400$ and linearly decaying to $t_{10}=200$. The learning rate for data fidelity optimization is $lr=100$.
Appendix D Additional Experiment Results and Implementation Details
Quantitative Results on FFHQ. Quantitative results for solving inverse problems on the FFHQ dataset are presented in Tables 10 and 11. Note that DCDP outperforms the baselines for various inverse problems.
Convergence Behavior. In Section 4.3 (Figure 8), we show that with a linear decaying time schedule for diffusion purification, DCDP iteratively improves the LPIPS scores and progressively converges towards a high-quality reconstruction. Similar convergence behavior holds when measuring the reconstruction quality using PSNR and SSIM, as shown in Figure 20. In this experiment, DCDP is applied to the nonlinear deblurring task, with the standard metrics averaged over 10 randomly selected images from the ImageNet dataset. The hyperparameters are the same as those reported in Appendix B (Table 8). Note that DCDP substantially improves upon the pure data fidelity optimization, implying diffusion purification is an efficient operator to enforce the image priors embedded in the pre-trained diffusion models.
Impact of Number of DDIM Steps. Here we provide experimental details for the ablation study on the impact of DDIM steps in Section 4.3. The experiments are conducted using a pixel space diffusion model based on DCDP (DCDP-DDIM) for solving the nonlinear deblurring problem. Except for the number of DDIM steps, all other hyperparameters are the same as those reported in Appendix B (Table 8). The standard metrics are averaged over 10 randomly selected images from the ImageNet dataset.
Impact of Number of Total Number of Iterations of $K$. Here we provide experimental details for the ablation study on the impact of $K$ in Section 4.3. The experiments are conducted using pixel space diffusion models (both DCDP-DDIM and DCDP-Tweedie). Except for $K$, all other hyperparameters are the same as those reported in Appendix B (Table 8). The standard metrics are averaged over 10 randomly selected images from the ImageNet dataset.
Figure 20: Convergence of our methods in terms of PSNR, SSIM and LPIPS for solving nonlinear deblurring. DCDP greatly improves upon the pure data fidelity optimization by leveraging the image priors embedded in the diffusion models.